OpenStudy (anonymous):

Number of Ni atoms in 3.4moles of Ni?

OpenStudy (accessdenied):

Do you know how many Ni atoms are in 1 mole of Ni? This would be the same as the number of things in 1 mole of anything.

OpenStudy (anonymous):

Do you mean grams?

OpenStudy (anonymous):

The [The problem doesn't say I would assume 6.022*10^23

OpenStudy (accessdenied):

1 mole of something will always be 6.022*10^23 somethings. That's just a standard set up for all mole measurements. It is kind of like speaking of dozens. 1 dozen of something is always 12 somethings. Or 3 dozens of something is 3*12 = 36 somethings. So, 1 mole of Ni is 6.022*10^23 Ni atoms. And 3 moles of Ni is 3*6.022*10^23 Ni atoms. Does that comparison make some sense?

OpenStudy (anonymous):

Yes, it does thank you!

OpenStudy (accessdenied):

Glad to help! That was the way I was introduced into moles, and I think it sets up a great intuition for what a mole really is, since it seems so obscure when you start learning about them. :) The more mathematical process is by unit conversions: x moles * (6.022*10^23 atoms / mole) = x*6.022*10^23 atoms. It's the same result as in the analogy above, but now we involved the units so you can see the conversion factor and unit cancellation. I find unit conversions great later on when you use molar mass (g/mol) to convert between grams and moles!